A ball bounces 1/3 of the height from which it is dropped. If it reaches a height of 3 units on the third bounce, from what height was it originally dropped ?
A 3 b 9 c 27 d 81 e 243


Answer by Alan3354(69443)

A ball bounces 1/3 of the height from which it is dropped. If it reaches a height of 3 units on the third bounce, from what height was it originally dropped ?
A 3 b 9 c 27 d 81 e 243
--------------
h = (1/3)(h0) = (1/3)^n*(h0)
h0 = h/(1/3)^n = 3/(1/3)^3 = 3*3^3
h0 = 81 units
-------------
b#...h
1.....27
2.....9
3.....3